OpenStudy (aizhalee):

Will Fan and Medal ! Kyra is using rectangular tiles of two types for a floor design. A tile of each type is shown below:

OpenStudy (aizhalee):

Which statement is correct?

OpenStudy (aizhalee):

I need help with understanding this please

OpenStudy (aizhalee):

@sdfgsdfgs can u help me with this one please

OpenStudy (aizhalee):

LOL omg 5 viewing but still no help D:

OpenStudy (anonymous):

you got it right though? what dont you understand?

OpenStudy (aizhalee):

I guessed on it :/ and got it right , but I want to know to do it correctly

OpenStudy (aizhalee):

Thank you for replying

OpenStudy (anonymous):

ok, so you know how the cordinates were (5,12) for S and (5'7) for P

OpenStudy (sdfgsdfgs):

read this plz - http://openstudy.com/study#/updates/529f8a70e4b01a67794e7625

OpenStudy (aizhalee):

Yes I see the points

OpenStudy (anonymous):

ok so what plus 7 gets you to 12?

OpenStudy (anonymous):

since that is the Y cordinate

OpenStudy (aizhalee):

5 +7 = 12

OpenStudy (anonymous):

!the last option is correct! why?

OpenStudy (anonymous):

ok so thats how you get 5

OpenStudy (aizhalee):

Ok whats next ?

OpenStudy (anonymous):

wait!

OpenStudy (aizhalee):

Thank you @sdfgsdfgs

OpenStudy (anonymous):

now you do the same for SR

OpenStudy (aizhalee):

@eninone B is correct,

OpenStudy (aizhalee):

Ok 12+0=2

OpenStudy (anonymous):

nooo

OpenStudy (anonymous):

cordinate S=what

OpenStudy (aizhalee):

5,12

OpenStudy (anonymous):

ok what is cordinate R?

OpenStudy (aizhalee):

9,12

OpenStudy (anonymous):

what plus 5 gives you 9

OpenStudy (aizhalee):

4

OpenStudy (anonymous):

so thats how you got the ratio 5:4

OpenStudy (anonymous):

do you get it now?

OpenStudy (aizhalee):

Yes I get it :D , so I just do it for MJ and ML

OpenStudy (aizhalee):

Thank you Im adding it in my notes

OpenStudy (anonymous):

yes, and that how you get the ratio for that too!!! good job and thanks for having pactients i didnt know how to explain it all in one piece

OpenStudy (aizhalee):

Thanks :) and you're very welcome

OpenStudy (anonymous):

\[similarity\ concepts\ says\ that \ two \ figures\ are\ similar\\ if \ the\ ratio\ of \ their\ corresponding\ sides\ are\ equal\ (constant).\\in\ the\ given \ fiqures \ only\ last \ statemet\ satisfy \ this \ condition.\\here,\ \frac{ SR }{ ML}=\frac{ 9-5 }{ 6-4}=\frac{ 4 }{ 2 }=\frac{ 2 }{ 1}\\ \frac{ PQ }{ JK }=\frac{ 9-5 }{ 6-4}=\frac{ 4 }{ 2 }=\frac{ 2 }{ 1 }\\ so\ last\ statement \ is\ correct!\]

OpenStudy (anonymous):

this solution is partially wrong. missed one line,. Option B is correct. overworked! time to go bed!